16) Most of the metal in the Model-T, a car built by Henry Ford from 1908 to 1926, was an alloy of steel and vanadium. Mixing a small amount of vanadium in with the steel makes it a little lighter, but it still maintains much of its strength. There are two isotopes of vanadium: V-50 (49.947 amu) and V-51 (50.944 amu). What is the percent abundance of each isotope if the average atomic mass of vanadium is 50.942? 17) When you bend a piece of Indium, it lets out a high-pitched shriek. There are two naturally occurring isotopes of Indium: In-113 (112.904 amu) and In-115 (114.904 amu). What is the percentage abundance of each isotope?
Answer #1

16)

average atomicmas = atomicmass of V-51*x + atomicmass of V-50*(100-x)/100

average atomicmas of v = 50.942 amu

atomicmass of V-51 = 50.944 amu

atomicmass of V-50 = 49.947 amu

50.942 = (50.944*x+49.947*(100-x))/100

percentage of V-51 = x = 99.8%

percentage of V-50 = 100-99.8 = 0.2%

17)

average atomicmas OF Indium is not given.please check it.